OpenStudy (anonymous):

What is the kinetic energy of a 1.0-kg billiard ball that moves at 5.0 m/s? A. 5.0 J B. 6.5 J C. 10.0 J D. 12.5 J

OpenStudy (aaronq):

Use the formula: \(\sf KE=mv^2\)

OpenStudy (anonymous):

.-. , doesnt make sence .-.

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

:{

OpenStudy (anonymous):

u left

OpenStudy (aaronq):

How does it not make sense?

OpenStudy (anonymous):

i dont know whatt to do

OpenStudy (aaronq):

oh crap i left out the 1/2 \(\sf KE=\dfrac{1}{2}mv^2\) m =mass v = velocity